Please help!

Prestige Worldwide - Trial balance as of December 31, 2018:

Gross Service Revenue $ 115,981,250

Bad Debt Expense $ 2,655,000

Discounts on Sales $ (3,575,000)

Salaries $ 55,825,000

Benefits $ 14,178,292

Administrative $ 4,342,500

Insurance $ 325,000

Interest $ 164,575

Depreciation $ 3,200,000

Maintenance $ 435,500

Purchased Services $ 6,146,500

Rent $ 2,240,000

Supplies $ 9,169,813

Utilities-Electricity $ 356,000

Cash and Cash Equivalents $ 9,545,000

Short-term Investments $ 2,750,000

Net Accounts Receivables $ 7,125,000

Inventory - Supplies $ 1,135,000

Property and Equipment $ 90,500,000

Accumulated Depreciation $ (43,500,000)

Investment Income $ 94,500

Account Payable $ 2,375,000

Accrued Expenses $ 9,125,000

Notes Payable $ 16,515,000

Contributions Received by Prestige $ 1,000,000

Long-Term Debt $ 26,500,000

Retained Earnings $ 13,040,000

Net revenues will reflect on Financial Statements what the finance organization actually expects to collect. Prestige Worldwide has a tax rate of 25% for 2018.

1. Using the information given above, construct the Income Statement and Balance Sheet for the Fiscal Year ended December 31, 2018.

2.If the benchmark for operating margin ratio is 20.06% for the entertainment industry, how does Prestige World wide’s compare to this benchmark? What are the implications to Prestige Worldwide of this ratio?

3. If the benchmark for net profit margin is 9.86% for the industry, how does PW's compare to this benchmark? What are the implications to PW's of this ratio?

4. What is the estimated Cash Flow using only the income statement or estimate the amount of cash generated during FY 2018 using the income statement. (Cash Flow = Net Income+Noncash expenses). The only Non-cash expense listed is depreciation expense. Why is this calculation Important to Manage?

5. What is PW’s Times Interest Earned ratio?
Answer #1

1.

Income Statement

Particulars $ Debit Particulars $ Credit
Purchased Services 6,146,500 Gross Service Revenue 115,981,250
Discounts on Sales 3,575,000 Investment Income 94,500
Bad Debt Expense 2,655,000
Salaries 55,825,000
Benefits 14,178,292
Administrative 4,342,500
Insurance 325,000
Interest 164,575
Depreciation 3,200,000
Maintenance 435,500
Rent 2,240,000
Supplies 9,169,813
Tax @25% 3454642.5
Net profit/income 10363927.5
Total 116075750 Total 116075750

Balance Sheet

Liabilities $ Assets $
Accumulated Depreciation (43,500,000) Utilities-Electricity 356,000
Account Payable 2,375,000 Cash and Cash Equivalents 9,545,000
Accrued Expenses 9,125,000 Short-term Investments 2,750,000
Notes Payable 16,515,000 Net Accounts Receivables 7,125,000
Contributions Received by Prestige 1,000,000 Inventory - Supplies 1,135,000
Long-Term Debt 26,500,000 Property and Equipment 90,500,000
Retained Earnings 13,040,000 Suspense (bal. fig.) 644000
Total 112055000 Total 112055000

2.

Operating margin = Operating profit / Net sales

15064645/ 112406250 = 13.4%

Operating margin of Prestige is lower than industry benchmark that is 20.06%, it implies prestige performing low.

3.

Net profit margin = net profit / net income

10363927.5 / 112500750 = 9.21%

Net profit margin of Prestige is almost equals to industry bench mark tat is 9.86%, it implites prestige performing at par with industry performance

4. Cash Flow = Net Income+Noncash expenses

= 10363927.5 + 3,200,000 = 13563927.5

5. Interest earned ratio = EBIT / Interest

= 13983145 / 164,575 = 84.96 times